OpenStudy (anonymous):

solve using the zero-factor property (t-4)(t-4)=0

OpenStudy (rajee_sam):

Zero-factor (or) Zero-Product property states that if a . b = 0 ; then a = 0 or b = 0 So if (t-4)(t-4) = 0 means ; t-4 = 0 ; t-4 = 0 S t = 4 and t=4 are the two solutions.